# Latent Wavelet Diffusion: Enabling 4K Image Synthesis for Free

### v2

@@ Abstract
High-resolution image synthesis remains a core challenge in generative modeling, particularly in balancing computational efficiency with the preservation of fine-grained visual detail. We present Latent Wavelet Diffusion (LWD), a lightweight framework that enables any latent diffusion model to scale to ultra-high-resolution image generation (2K to 4K) for free. LWD introduces three key components: (1) a scale-consistent variational autoencoder objective that enhances the spectral fidelity of latent representations; (2) wavelet energy maps that identify and localize detail-rich spatial regions within the latent space; and (3) a time-dependent masking strategy that focuses denoising supervision on high-frequency components during training. LWD requires no architectural modifications and incurs no additional computational overhead. Despite its simplicity, it consistently improves perceptual quality and reduces FID in ultra-high-resolution image synthesis, outperforming strong baseline models. These results highlight the effectiveness of frequency-aware, signal-driven supervision as a principled and efficient approach for high-resolution generative modeling.
@@ 1 Introduction
Diffusion models have become a dominant paradigm in generative modeling, achieving state-of-the-art results in tasks such as image synthesis podell2024sdxl ; sauer2024adversarial ; gwit ; wang2025designdiffusionhighqualitytexttodesignimage , super-resolution saharia2022image ; shipin ; wang2024sinsr , and image editing brooks2023instructpix2pix ; meng2022sdedit . Foundational methods like Denoising Diffusion Probabilistic Models (DDPM) ho2020denoising and Denoising Diffusion Implicit Models (DDIM) song2021denoising have enabled increasingly powerful variants. Latent Diffusion Models (LDMs) rombach2022high improve efficiency by operating in a learned latent space, while recent architectures such as Diffusion Transformers (DiTs) peebles2023scalable ; esser2024scaling further enhance scalability and modeling capacity.
Despite recent advances, generating ultra-high-resolution (UHR) images at resolutions of 2K to 4K and beyond remains a significant challenge. Simply scaling models trained on lower-resolution data often fails to generalize, leading to repeated structures, blurred textures, and spatial inconsistencies bar2023multidiffusion ; he2023scalecrafter . Naive alternatives, such as hierarchical generation pipelines and post-hoc super-resolution techniques du2024demofusion ; he2023scalecrafter , typically produce oversmoothed outputs. To overcome these limitations, several strategies have been explored. One direction involves direct UHR training or fine-tuning ren2024ultrapixel ; pixartsigma , although this typically demands extensive computational resources, access to proprietary high-resolution datasets 2024SoraReview , and substantial GPU memory for large model backbones flux ; esser2024scaling . Other efforts focus on architectural modifications to improve long-range dependency modeling liu2024unified , or optimization techniques that enhance the quality of latent representations pmlr-v235-hahm24a .
@@ 2 Related Work
Diffusion Models and Latent Generation. Diffusion models have become foundational in generative modeling, particularly for image synthesis anonymous2025efficient ; condimgsyntsurvey , by progressively denoising Gaussian noise using a learned score function. Variants based on stochastic differential equations song2021scorebased , probability flow ordinary differential equations lipman2023flow , and reinforcement-trained objectives black2024training have expanded the design space with improved stability and sampling efficiency.
@@
To reduce the cost of high-resolution generation, Latent Diffusion Models (LDMs) rombach2022high perform synthesis in a compressed space learned via variational autoencoders (VAEs). However, generation quality is closely tied to the spectral fidelity and structure of these latent representations. Prior work has sought to improve this through enhanced VAE architectures esser2021taming , hierarchical compression takida2024hqvae , and frequency-aware regularization improvediffusability ; kouzelis2025eq . We build on this direction by integrating frequency-sensitive supervision both during encoding and throughout the denoising process.
@@
Flow Matching and High-Resolution Diffusion. Flow matching lipman2023flow ; esser2024scaling offers an alternative to classical diffusion by learning a continuous velocity field that maps noise to data in latent space, eliminating the need for fixed noise schedules. This formulation underlies models such as Flux flux , which, paired with DiT backbones, has demonstrated strong performance in ultra-high-resolution pipelines diffusion4k ; urae . Our method extends this family by introducing frequency-based spatial masking into the flow-matching objective. Through wavelet decomposition of the latent space, LWD computes spatial saliency maps that guide targeted supervision toward detail-rich regions, enhancing fine structure without modifying model architecture.
@@
Variational Autoencoder Optimization. The performance of latent diffusion models at high resolutions depends critically on the expressiveness and spectral consistency of the VAE. Improvements include multi-scale encoders vahdat2020nvae ; takida2024hqvae , spectral loss functions 9746027 , and scale-consistency constraints improvediffusability ; kouzelis2025eq . Wavelet-based methods esteves2024spectral ; lin2024open ; agarwal2025cosmos enrich latent expressiveness by isolating frequency components, while compression-oriented approaches xie2025sana ; tang2024vidtok ; hacohen2024ltx aim to reduce token count for improved sampling efficiency. Our LWD fine-tunes a pretrained VAE with a scale-consistent spectral loss that suppresses spurious high-frequency noise. This regularized latent space facilitates downstream wavelet decomposition and supports our spatially adaptive denoising objective.
@@
Ultra High-Resolution Image Synthesis. Maintaining global structure and fine detail in ultra-resolution generation is a persistent challenge. Standard diffusion models tend to produce repetitive patterns or distortions he2023scalecrafter . Existing solutions include cascaded generation ho2022cascaded , progressive upsampling gu2024matryoshka , domain-specific pipelines ren2024ultrapixel , and latent-space super-resolution jeong2025latent . However, training-based methods xie2023difffit ; zheng2024any ; guo2024make ; pixartsigma are resource-intensive, and training-free approaches bar2023multidiffusion ; lee2023syncdiffusion often yield local artifacts.
@@
Methods such as ScaleCrafter he2023scalecrafter mitigate repetition through dilated convolutions but may distort structure, while ResMaster shi2025resmaster uses low-resolution references for guided refinement. HiDiffusion zhang2023adding introduces architectural changes that risk performance trade-offs, and progressive strategies like DemoFusion du2024demofusion suffer from slow inference and irregular patterns. Diffusion-4K diffusion4k and URAE urae advance latent modeling and parameter-efficient adaptation at 4K resolution. Our LWD complements these approaches by introducing signal-driven, spatially adaptive supervision that improves structural and perceptual fidelity for free.
@@
Generative Modeling in the Frequency Domain. Frequency structure plays an increasingly important role in generative modeling diffmadeslim . Wavelet-based diffusion methods, such as WaveDiff phung2023wavelet , and spectral decomposition approaches have proven useful for efficient sampling qian2024boosting , super-resolution aloisi2024wavelet ; sigillo2025quaternionwaveletdiff , and restoration wavedm ; zhao2024wavelet ; lowlight . FouriScale huang2024fouriscale demonstrated that frequency filtering enhances coherence, and DiffuseHigh kim2025diffusehigh leveraged low-frequency DWT guidance to improve global structure in UHR synthesis.
@@
Diffusion-4K diffusion4k incorporated wavelet losses in the latent space to balance frequency bands, but applied them uniformly across all spatial locations. In contrast, LWD introduces wavelet-based spatial conditioning through time-dependent masking. Rather than treating frequency as a passive loss signal, our LWD actively uses local frequency energy to modulate supervision across space and time, concentrating on learning where detail matters most. This enables sharper, more coherent synthesis without modifying the underlying model or increasing inference cost.
@@ 3 Methodology
LWD comprises two sequential stages. First, we fine-tune a variational autoencoder (VAE) using a scale-consistency objective to obtain spectrally stable and semantically meaningful latent representations. This step is performed independently and prepares the latent space for subsequent frequency-based modulation.
@@ 3.1 VAE Fine-Tuning with Scale-Consistency Loss
Here, we incorporate a scale-consistency term improvediffusability ; kouzelis2025eq that encourages the VAE to maintain structural coherence across resolution scales. This regularization naturally suppresses high-frequency artifacts while ensuring robust encoding of essential structural information in $z$, which is critical for the downstream wavelet decomposition in LWD.
Unlike recent approaches that inject frequency conditioning directly into the encoder aloisi2024wavelet ; diffusion4k , our formulation decouples signal regularization and generation: we first sculpt the latent space to exhibit desirable frequency properties, and then use this structure to guide the training of the diffusion model. This preserves architectural modularity while enabling effective frequency-aware supervision.
@@ 3.2 Wavelet-Derived Frequency Saliency Maps
This map serves as a proxy for local structural richness, highlighting regions of the latent space associated with high-frequency content (e.g., textures, contours, transitions). Unlike learned attention mechanisms based on semantic similarity (e.g., DINO caron2021emerging ), our approach is deterministic and directly derived from signal properties. While we refer to $A_{\text{wavelet}}$ as an "attention map" for interpretability, it is best understood as a frequency-aware saliency measure.
@@ 3.3 Adaptive Flow Matching with Frequency-Guided Masking
We adopt a continuous-time flow-matching formulation (lipman2023flow, ; esser2024scaling, ) for training the latent diffusion model. Given a target latent $z_{0}$ and noise sample $\epsilon\sim\mathcal{N}(0,I)$, we define interpolated samples as:
@@
While selective spatial supervision has been investigated using transformer attention yoda , our wavelet-derived saliency offers a fundamentally different perspective based on signal processing principles rather than learned semantic features, offering computational advantages and a more generalizable solution.
@@ 4.1 Experimental Setup
Datasets. We evaluate LWD on two ultra-resolution datasets. Aesthetic-4K (diffusion4k, ) is a curated 4K benchmark with GPT-4o-generated captions and high visual quality. LAION-High-Res is a filtered subset of LAION-5B (schuhmann2022laion, ), from which we sample 50K 2K-resolution and 20K 4K-resolution image-caption pairs. These two datasets differ in both visual and linguistic distributions, allowing us to assess both generation fidelity and generalization under caption variance.
@@
Implementation Details. We implement LWD using PyTorch and pytorch-wavelets (cotter_2019, ) for Haar-based DWT. VAE fine-tuning is performed for 60K steps with a batch size of 4 and a learning rate of $1\times 10^{-5}$ on NVIDIA A100 GPUs. The frequency-guided loss is applied during the training of the architectures we tested. We set the masking lower bound $\ell=0.3$, selected via ablation to ensure each spatial location receives at least 30% of refinement steps. We will release the code on GitHub after the review process. Training requirements varied across architectures. LWD+URAE (Flux-based) flux ; urae required approximately 4 hours and 2k iterations at batch size 1 for the 2048 version, while the 4096 version needed 24 hours for 2k iterations, significantly fewer than the original work, which suggests 10k iterations. Diffusion4K (SD3-based) esser2024scaling ; diffusion4k was trained for 48 hours at 10k iterations (half the originally required iterations) with batch size 8 for the 2048 version, maintaining identical parameters for the 4096 version except with batch size 4. SANA xie2025sana models (both 2048 and 4096 versions) were trained for 24 hours, achieving 33k iterations (one-third of the originally suggested iterations) with batch sizes of 2 and 1, respectively. PixArt-$\Sigma$ pixartsigma required only 24 hours and 1.5k iterations at batch size 2, compared to the suggested 15k iterations.
Evaluation Metrics. To comprehensively evaluate ultra-resolution text-to-image generation, we report multiple complementary metrics across three dimensions: image quality (Fréchet Inception Distance (FID), LPIPS), semantic alignment (CLIPScore (hessel2021clipscore, ), QualiCLIP (agnolucci2024qualityaware, )), and perceptual quality (MAN-IQA (yang2022maniqa, ), HPSv2.1 (wu2023humanpreferencescorev2, ), PickScore (kirstain2023pick, )). Note that lower values indicate better performance for FID and LPIPS, while higher values are preferred for all other metrics. We assess texture fidelity with the Gray Level Co-occurrence Matrix (GLCM) Score diffusion4k and fine-grained detail preservation using JPEG Compression Ratio to measure high-frequency compressibility as a perceptual quality proxy.
Table 1: Quantitative results on different metrics. The prompts are from the HPD wu2023humanpreferencescorev2 dataset. All images are at a resolution of 2048 $\times$ 2048.
@@
SDEdit meng2022sdedit 35.59 0.6456 0.3736 0.4480 30.92 22.86
I-Max du2024imaxmaximizeresolutionpotential 36.28 0.6750 0.3641 0.4139 30.62 23.02
PixArt-Sigma-XL pixartsigma 36.58 0.6801 0.2949 0.4438 30.66 22.92
Sana-1.6B xie2025sana 35.75 0.7169 0.3666 0.5796 30.42 22.83
Lumina-Image 2.0 qin2025lumina 54.96 0.6445 0.3663 0.4567 23.08 21.15
FLUX-1.dev flux 37.58 0.6371 0.4110 0.5468 28.73 22.68
URAE urae 35.25 0.6717 0.4076 0.5423 31.15 22.41

### v3

# Latent Wavelet Diffusion For Ultra-High-Resolution Image Synthesis

@@ Abstract
High-resolution image synthesis remains a core challenge in generative modeling, particularly in balancing computational efficiency with the preservation of fine-grained visual detail. We present Latent Wavelet Diffusion (LWD), a lightweight training framework that significantly improves detail and texture fidelity in ultra-high-resolution (2K-4K) image synthesis. LWD introduces a novel, frequency-aware masking strategy derived from wavelet energy maps, which dynamically focuses the training process on detail-rich regions of the latent space. This is complemented by a scale-consistent VAE objective to ensure high spectral fidelity. The primary advantage of our approach is its efficiency: LWD requires no architectural modifications and adds zero additional cost during inference, making it a practical solution for scaling existing models. Across multiple strong baselines, LWD consistently improves perceptual quality and FID scores, demonstrating the power of signal-driven supervision as a principled and efficient path toward high-resolution generative modeling.
@@ 1 Introduction
Diffusion models have become a dominant paradigm in generative modeling, achieving state-of-the-art results in tasks such as image synthesis (Podell et al., 2024; Sauer et al., 2024; Lopez et al., 2025; Wang et al., 2025), super-resolution (Saharia et al., 2022; Sigillo et al., 2024; Wang et al., 2024), and image editing (Brooks et al., 2023; Meng et al., 2022). Foundational methods like Denoising Diffusion Probabilistic Models (DDPM) Ho et al. (2020) and Denoising Diffusion Implicit Models (DDIM) Song et al. (2021a) have enabled increasingly powerful variants. Latent Diffusion Models (LDMs) Rombach et al. (2022) improve efficiency by operating in a learned latent space, while recent architectures such as Diffusion Transformers (DiTs) (Peebles & Xie, 2023; Esser et al., 2024) further enhance scalability and modeling capacity.
Despite recent advances, generating ultra-high-resolution (UHR) images at resolutions of 2K to 4K and beyond remains a significant challenge. Simply scaling models trained on lower-resolution data often fails to generalize, leading to repeated structures, blurred textures, and spatial inconsistencies (Bar-Tal et al., 2023; He et al., 2023). Naive alternatives, such as hierarchical generation pipelines and post-hoc super-resolution techniques (Du et al., 2024a; He et al., 2023), typically produce oversmoothed outputs. To overcome these limitations, several strategies have been explored. One direction involves direct UHR training or fine-tuning (Ren et al., 2024; Chen et al., 2024), although this typically demands extensive computational resources, access to proprietary high-resolution datasets (Liu et al., 2024b), and substantial GPU memory for large model backbones (Labs, 2024; Esser et al., 2024). Other efforts focus on architectural modifications to improve long-range dependency modeling (Liu et al., 2024a), or optimization techniques that enhance the quality of latent representations (Hahm et al., 2024).
@@ 2 Related Work
Diffusion Models and Latent Generation. Diffusion models have become foundational in generative modeling, particularly for image synthesis (Shen et al., 2025; Zhan et al., 2024), by progressively denoising Gaussian noise using a learned score function. Variants based on stochastic differential equations (Song et al., 2021b), probability flow ordinary differential equations (Lipman et al., 2023), and reinforcement-trained objectives (Black et al., 2024) have expanded the design space with improved stability and sampling efficiency.
@@
To reduce the cost of high-resolution generation, Latent Diffusion Models (LDMs) (Rombach et al., 2022) perform synthesis in a compressed space learned via variational autoencoders (VAEs). However, generation quality is closely tied to the spectral fidelity and structure of these latent representations. Prior work has sought to improve this through enhanced VAE architectures (Esser et al., 2021), hierarchical compression (Takida et al., 2024), and frequency-aware regularization (Skorokhodov et al., 2025; Kouzelis et al., 2025). We build on this direction by integrating frequency-sensitive supervision both during encoding and throughout the denoising process.
@@
Flow Matching and High-Resolution Diffusion. Flow matching Lipman et al. (2023); Esser et al. (2024) offers an alternative to classical diffusion by learning a continuous velocity field that maps noise to data in latent space, eliminating the need for fixed noise schedules. This formulation underlies models such as Flux Labs (2024), which, paired with DiT backbones, has demonstrated strong performance in ultra-high-resolution pipelines (Zhang et al., 2025; Yu et al., 2025). Our method extends this family by introducing frequency-based spatial masking into the flow-matching objective. Through wavelet decomposition of the latent space, LWD computes spatial saliency maps that guide targeted supervision toward detail-rich regions, enhancing fine structure without modifying model architecture.
@@
Variational Autoencoder Optimization. The performance of latent diffusion models at high resolutions depends critically on the expressiveness and spectral consistency of the VAE. Improvements include multi-scale encoders (Vahdat & Kautz, 2020; Takida et al., 2024), spectral loss functions (Björk et al., 2022), and scale-consistency constraints (Skorokhodov et al., 2025; Kouzelis et al., 2025). Wavelet-based methods (Esteves et al., 2024; Lin et al., 2024; Agarwal et al., 2025) enrich latent expressiveness by isolating frequency components, while compression-oriented approaches (Xie et al., 2025; Tang et al., 2024; HaCohen et al., 2024) aim to reduce token count for improved sampling efficiency. Our LWD fine-tunes a pretrained VAE with a scale-consistent spectral loss that suppresses spurious high-frequency noise. This regularized latent space facilitates downstream wavelet decomposition and supports our spatially adaptive denoising objective.
@@
Ultra High-Resolution Image Synthesis. Maintaining global structure and fine detail in ultra-resolution generation is a persistent challenge. Standard diffusion models tend to produce repetitive patterns or distortions (He et al., 2023). Existing solutions include cascaded generation (Ho et al., 2022), progressive upsampling (Gu et al., 2024), domain-specific pipelines (Ren et al., 2024), and latent-space super-resolution (Jeong et al., 2025). However, training-based methods (Xie et al., 2023; Zheng et al., 2024; Guo et al., 2024; Chen et al., 2024) are resource-intensive, and training-free approaches (Bar-Tal et al., 2023; Lee et al., 2023) often yield local artifacts.
@@
Methods such as ScaleCrafter He et al. (2023) mitigate repetition through dilated convolutions but may distort structure, while ResMaster Shi et al. (2025) uses low-resolution references for guided refinement. HiDiffusion Zhang et al. (2023) introduces architectural changes that risk performance trade-offs, and progressive strategies like DemoFusion Du et al. (2024a) suffer from slow inference and irregular patterns. Diffusion-4K Zhang et al. (2025) and URAE Yu et al. (2025) advance latent modeling and parameter-efficient adaptation at 4K resolution. Our LWD complements these approaches by introducing signal-driven, spatially adaptive supervision, which improves structural and perceptual fidelity at no additional cost.
@@
Generative Modeling in the Frequency Domain. Frequency structure plays an increasingly important role in generative modeling (Yang et al., 2023). Wavelet-based diffusion methods, such as WaveDiff Phung et al. (2023), and spectral decomposition approaches have proven useful for efficient sampling (Qian et al., 2024), super-resolution (Aloisi et al., 2024; Sigillo et al., 2025), and restoration (Huang et al., 2024b; Zhao et al., 2024; Jiang et al., 2023). FouriScale Huang et al. (2024a) demonstrated that frequency filtering enhances coherence, and DiffuseHigh Kim et al. (2025) leveraged low-frequency DWT guidance to improve global structure in UHR synthesis.
@@
Diffusion-4K (Zhang et al., 2025) incorporated wavelet losses in the latent space to balance frequency bands, but applied them uniformly across all spatial locations. In contrast, LWD introduces wavelet-based spatial conditioning through time-dependent masking. Rather than treating frequency as a passive loss signal, our LWD actively uses local frequency energy to modulate supervision across space and time, concentrating on learning where detail matters most. This enables sharper, more coherent synthesis without modifying the underlying model or increasing inference cost.
@@ 3 Methodology
LWD operates in two sequential stages. The first stage fine-tunes a variational autoencoder (VAE) using a scale-consistency objective. This independent step yields spectrally stable latent representations, preparing the space for the subsequent frequency-based modulation.
@@ 3.1 VAE Fine-Tuning with Scale-Consistency Loss
Here, we incorporate a scale-consistency term (Skorokhodov et al., 2025; Kouzelis et al., 2025) that encourages the VAE to maintain structural coherence across resolution scales. This regularization naturally suppresses high-frequency artifacts while ensuring robust encoding of essential structural information in $z$, which is critical for the downstream wavelet decomposition in LWD.
Unlike recent approaches that inject frequency conditioning directly into the encoder (Aloisi et al., 2024; Zhang et al., 2025), our formulation decouples signal regularization and generation: we first sculpt the latent space to exhibit desirable frequency properties, and then use this structure to guide the training of the diffusion model. This preserves architectural modularity while enabling effective frequency-aware supervision.
@@ 3.2 Wavelet-Derived Frequency Saliency Maps
This map serves as a proxy for local structural richness, highlighting regions of the latent space associated with high-frequency content (e.g., textures, contours, transitions). Unlike learned attention mechanisms based on semantic similarity (e.g., DINO (Caron et al., 2021)), our approach is deterministic and directly derived from signal properties. While we refer to $A_{\text{wavelet}}$ as an ”attention map” for interpretability, it is best understood as a frequency-aware saliency measure.
@@ 3.3 Adaptive Flow Matching with Frequency-Guided Masking
We adopt a continuous-time flow-matching formulation (Lipman et al., 2023; Esser et al., 2024) for training the latent diffusion model. Given a target latent $z_{0}$ and noise sample $\epsilon\sim\mathcal{N}(0,I)$, we define interpolated samples as:
@@
While selective spatial supervision has been investigated using transformer attention (Moser et al., 2025), our wavelet-derived saliency offers a fundamentally different perspective based on signal processing principles rather than learned semantic features, offering computational advantages and a more generalizable solution.
@@ 4.1 Experimental Setup
Datasets. We evaluate LWD on two ultra-resolution datasets. Aesthetic-4K Zhang et al. (2025) is a curated 4K benchmark with GPT-4o-generated captions and high visual quality. LAION-High-Res is a filtered subset of LAION-5B Schuhmann et al. (2022), from which we sample 50K 2K-resolution and 20K 4K-resolution image-caption pairs. These two datasets differ in both visual and linguistic distributions, allowing us to assess both generation fidelity and generalization under caption variance.
@@
Implementation Details. We implement LWD using PyTorch and pytorch-wavelets Cotter (2019) for Haar-based DWT. VAE fine-tuning is performed for 60K steps with a batch size of 4 and a learning rate of $1\times 10^{-5}$ on NVIDIA A100 GPUs. The frequency-guided loss is applied during the training of the architectures we tested. We set the masking lower bound $\ell=0.3$, selected via ablation to ensure each spatial location receives at least 30% of refinement steps. We will release the code on GitHub after the review process. Training requirements varied across architectures. LWD+URAE (Flux-based) Labs (2024); Yu et al. (2025) required approximately 4 hours and 2k iterations at batch size 1 for the 2048 version, while the 4096 version needed 24 hours for 2k iterations, significantly fewer than the original work, which suggests 10k iterations. Diffusion4K (SD3-based) (Esser et al., 2024; Zhang et al., 2025) was trained for 48 hours at 10k iterations (half the originally required iterations) with batch size 8 for the 2048 version, maintaining identical parameters for the 4096 version except with batch size 4. SANA Xie et al. (2025) models (both 2048 and 4096 versions) were trained for 24 hours, achieving 33k iterations (one-third of the originally suggested iterations) with batch sizes of 2 and 1, respectively. PixArt-$\Sigma$ Chen et al. (2024) required only 24 hours and 1.5k iterations at batch size 2, compared to the suggested 15k iterations.
Evaluation Metrics. We evaluate ultra-resolution text-to-image generation across three key dimensions. For image quality, we use Fréchet Inception Distance (FID) and LPIPS (lower is better), alongside the Gray Level Co-occurrence Matrix (GLCM) Score for texture and JPEG Compression Ratio as a proxy for fine-grained detail. For semantic alignment, we report CLIPScore (Hessel et al., 2021) and QualiCLIP Agnolucci et al. (2024). Finally, for perceptual quality, we use MAN-IQA Yang et al. (2022), HPSv2.1 Wu et al. (2023), and PickScore Kirstain et al. (2023). Higher values indicate better performance for all metrics except FID and LPIPS.
Table 1: Quantitative results on different metrics. The prompts are from the HPD (Wu et al., 2023) dataset. All images are at a resolution of 2048 $\times$ 2048.
@@
SDEdit Meng et al. (2022) 35.59 0.6456 0.3736 0.4480 30.92 22.86
I-Max Du et al. (2024b) 36.28 0.6750 0.3641 0.4139 30.62 23.02
Diffusion-4K Zhang et al. (2025) 37.10 0.6920 0.3550 0.4815 30.55 22.80
PixArt-Sigma-XL Chen et al. (2024) 36.58 0.6801 0.2949 0.4438 30.66 22.92
Sana-1.6B Xie et al. (2025) 35.75 0.7169 0.3666 0.5796 30.42 22.83
Lumina-Image 2.0 Qin et al. (2025) 54.96 0.6445 0.3663 0.4567 23.08 21.15
FLUX-1.dev Labs (2024) 37.58 0.6371 0.4110 0.5468 28.73 22.68
URAE Yu et al. (2025) 35.25 0.6717 0.4076 0.5423 31.15 22.41